# Towards LifeSpan Cognitive Systems

### v1

@@ Abstract
Building a human-like system that continuously interacts with complex environments—whether simulated digital worlds or human society—presents several key challenges. Central to this is enabling continuous, high-frequency interactions, where the interactions are termed experiences. We refer to this envisioned system as the LifeSpan Cognitive System (LSCS). A critical feature of LSCS is its ability to engage in incremental and rapid updates while retaining and accurately recalling past experiences. We identify two major challenges in achieving this: (1) Abstraction and Experience Merging, and (2) Long-term Retention with Accurate Recalling. These properties are essential for storing new experiences, organizing past experiences, and responding to the environment in ways that leverage relevant historical data. Unlike language models with continual learning, which typically rely on large corpora for fine-tuning and focus on improving performance within specific domains or tasks, LSCS must rapidly and incrementally update with new information from its environment at a high frequency. Existing technologies with the potential of solving the above two major challenges can be classified into four classes based on a conceptual metric called Storage Complexity, which measures the relative space required to store past experiences. Each of these four classes of technologies has its own strengths and limitations. Given that none of the existing technologies can achieve LSCS alone, we propose a novel paradigm for LSCS that integrates all four classes of technologies. The new paradigm operates through two core processes: Absorbing Experiences and Generating Responses.
@@ 1 Introduction
Imagine an environment such as a simulated digital world [77], a virtual world like Minecraft [109], the Marvel Universe depicted in movies, or even the complexities of human society. Developing an intelligent system capable of engaging with such environments—interacting with its surroundings, absorbing information from experiences, self-evolving based on feedback [142], and living through an entire lifespan, either as a real or virtual entity—remains a significant challenge. When the system inhabits the environment, it interacts with everything around it, including humans, objects, and other elements of the environment. We define these interactions as the system’s experiences, encompassing all of the system’s behaviors and the feedback it receives from the environment—whether visual, physical, linguistic, or across any other modality. This system’s cognitive capabilities include the ability to perceive and interpret its surroundings, retain and recall information through memory, learn and adapt from feedback, and engage in reasoning and decision-making processes. These cognitive functions enable the system to navigate its environment intelligently and evolve over time. We denote the system with the above capabilities as the LifeSpan Cognitive System (LSCS).
@@
Table 1: In this paper, we discuss four categories of technologies with the sub-categories being sub-sections. Here “Abs & ExMg” refers to“Abstraction and Experiences Merging”, "Re & AcRe" means “Long-term Retention and Accurate Recalling”, and “SC” represents “Storage Complexity”.
@@
To achieve LSCS, some major challenges must be addressed: (1) Abstraction and Experience Merging: The system must distill experiences from its environment by extracting essential information and integrating these abstracted experiences with its existing memories. Interactions with the environment—whether through conversations [28, 29] or visual inputs [118]—often contain redundant information that should be filtered out before storing in memory. Once the key information is abstracted, this crucial information can be used to update the system, merging new abstracted experiences with previous ones while resolving potential conflicts. This process enables the system to acquire new skills, deepen its understanding, correct misconceptions, and continually evolve. (2) Long-term Retention and Accurate Recalling: For systems that store past experiences in various forms—whether in memory, context, knowledge bases, or model weights—it is crucial to accurately recall relevant information in response to current queries from the environment. This capability necessitates the model’s ability to remember events from the distant past and make informed decisions based on all past experiences. The above two properties distinguish constructing LSCS from existing challenges such as long context problems [113] or continual learning [121], where long context problems focus on solving a problem with extremely abundant information and continual learning mainly focuses on enabling models to capture ever-changing world knowledge [139, 47], facilitating self-evolution through human feedback [102], or adapting to specific domains [56, 125].
@@
There are existing technologies that address some of the challenges described above, but each typically falls short of the other. We define Storage Complexity as a conceptual measure of the storage requirements beyond model parameters, expressed as a function of past experiences (denoted as $\mathcal{E}$), such as the number of tokens from interactions between the system and the environment. These technologies can be categorized into four groups (illustrated in Figure 1), which will be elaborated in the subsequent sections: Saving $\mathcal{E}$ into Model Parameters (Section 2), Saving $\mathcal{E}$ into Explicit Memory (Section 3), Saving $\mathcal{E}$ into Knowledge Bases for Retrieval (Section 4), and Saving $\mathcal{E}$ into Raw Text and Process All (Section 5). The advantages and potential drawbacks of each approach are discussed in Section 6, with a high-level summary provided in Table 1. This table also outlines the structure of the paper, further subdividing some categories.
@@
Given the limitations of existing technologies, we propose a new formulation for LSCS, which integrates all the aforementioned approaches. The operational design is depicted in Figure 2, where the process is split into two stages: (1) Abstraction and Merging Experiences when new experiences arise. (2) Generating Responses based on queries from the environment, ensuring long-term retention and precise recall. Each component in this design corresponds to one of the technology categories discussed earlier, and their roles are explored in more detail in Section 7.
@@
Figure 2: The Operating Diagram of LSCS mainly include two parts: (1) Abstraction and Experiences Merging; (2) Generating responses according to the environmental query, where the ability of long-term retention and accurate recalling should be guaranteed.
@@ 2 Methods of Saving $\mathcal{E}$ into Model Parameters
Absorbing $\mathcal{E}$ via Model Editing [128]. Given past experiences, it is possible to construct a knowledge graph from these experiences and then convert the triplets in the graph into factual statements, which could be further injected into the model parameters via model editing. Here we focus on the methods that directly make edits on the model parameters, i.e., the new knowledge is saved in parameters. Methods such as ROME [72], MEMIT [73] propose a closed-form solution to edit the MLP layers, while Model-Editing-FT [30] fine-tunes the whole model on the factual statements that need to be injected into the model. T-Patcher [44] and CaliNet [26] propose to store the new knowledge in additional neurons, however, this may encounter the issue of ever-growing parameters in our setting. As model-editing methods typically focus on storing simple factual statements in the model, they are not directly applicable in the context of LSCS. Thus adjustments are needed for the application.
@@
Absorbing $\mathcal{E}$ via Continual Learning. Continual learning is usually used in the following situations: (1) Performing real-time assimilation of dynamic data. Existing methods can incorporate information from various sources such as news [98, 129], scholarly articles [19], and social media [19]. [98] presents ERNIE 2.0, which is a continual pre-training framework that incrementally builds and learns from multiple tasks to maximize knowledge extraction from training data. (2) Injecting a large amount of knowledge into the language model. [48] introduces continual knowledge learning, a method for updating temporal knowledge in LLMs, reducing forgetting while acquiring new information. (3) Domain adaptation. [19] continually trains the model on new data streams for both language and vision, and [56] introduces a soft-masking mechanism to update language models with domain corpora, aiming to boost performance while preserving general knowledge. Various models are proposed as the domain-adapted version of base language models such as financial domain [124], E-commerce domain [70], and academic content [119]. These techniques could all be used for LSCS to adjust to the most up-to-date information. (4) Teaching the system how to speak and how to perform certain types of reasoning. To enable the models to solve novel tasks, task-incremental continual instruction tuning is proposed [114, 112]. These tasks could be related to mathematical problems [4], calculators, search engines, and databases [83]. With the rapid emergence of new tools like advanced software libraries, novel APIs, or domain-specific utilities [64, 53], there is a growing need to continually update language models so they can quickly adapt and master these new tools. (5) Adapting to evolving societal values. This leads to continual alignment, where two categories exist: (i) Updating LLMs to align to changing societal values; (ii) Incorporating new demographic groups or value systems into LLMs. Some representative works include CPPO [134], COPF [132], COPR [133], which aim to prevent the forgetting of old policies while injecting incremental preferences objectives into the model. For LSCS, these techniques are important as the system may encounter situations where it needs to be up-to-date, learn plenty of new knowledge, adapt to a new domain if the system enters a new environment, learn a specific kind of tasks, or has to update to meet the societal values.
@@
Catastrophic Forgetting. One major limitation of Saving $\mathcal{E}$ into model parameters is catastrophic forgetting, which is a general problem for methods that require updating model parameters both in model-editing (e.g. in sequential editing scenarios, only LLMs with an external module can work in sequential editing tasks such as WISE [110], GRACE [38] while methods without an external module typically fail [72, 73]) and continual learning. LLMs trained on different stages can encounter unconscious forgetting [66], eroding their general knowledge through continual instruction tuning. Studies [82] have shown that the behavior of safely aligned LLMs can degrade under these conditions. Some metrics are proposed in TRACE [112] to measure the forgetting of LLMs. While we want to prevent catastrophic forgetting, we do intend to have some extent of forgetting, which is aligned with the ability of Abstraction and Experiences Merging. Different from unlearning part of the knowledge in the LLM which is conducted after having the undesired knowledge in the model [117], we focus on encouraging the model to abstract the information before injecting new experiences into model parameters, where it is acceptable to discard some details. However, Abstracting experiences and discarding unimportant details are not fully explored in continual learning. It is well known that a language model cannot fully remember the training data and instead extracts inherent statistical patterns from the dataset. This aligns closely with our goals in LSCS, and having better control over which information is discarded could yield even more benefits.
@@ 3.1 Methods with a fixed-sized memory module
Most existing works with fixed-sized memory modules have a small memory. Before the transformer era, plenty of works argument recurrent neural networks (RNNs) with a small memory module such as Memory Network [120] and its follow-up [97]. More advanced methods include TARDIS [33], ARMIN [63], Fast Weights [5], Ordered Memory [92], etc. Borrowing ideas from human memory, dual-memory system to mimic short-term and long-term memories is proposed [6, 7]. Later with the introduction of transformers, Memory Transformer [11] adds memory tokens at the beginning of the sequence to summarize the given sequence, while RMT [9, 10] proposes to add memory tokens both at the beginning and the end of the sequence. Here  [10] scales the number of memory tokens to 512 which enables long-term retention of previous information. EMMA [74] has a memory pool of similar size, including both short-term and long-term memory. Then Larimar [22] proposes to use a fixed-sized episodic memory which is shown to be capable of storing up to 1000 factual statements. As opposed to small fixed-sized memory, MemoryLLM [116] introduces a memory for Llama2-7B [104] encapsulating 1B parameters with 7680 memory tokens per layer, enormously enlarging the memory pool. Differently, Infini-Attention [75] proposes to use linear attention to attend to the past seen tokens, where the stored matrix performs exactly like a fixed-sized memory (of size $d_{key}\times(d_{value}+1)\times H\times L$, where $d_{key},d_{value},H,L$ refer to the dimension of key states, dimension of value states, number of heads and number of layers, respectively) abstracting the past tokens. These fixed-sized memory modules typically come with higher compression than flexible-sized memory modules and also requires much more training for deployment.
@@ 3.2 Methods with flexible-sized memory pool
Memory Module in Hidden Space. Some methods introduce memory slots alongside the inputs to encode information, where the number of memory slots vary according to the length of the input document [2]. Other techniques store key-value pairs in a memory pool for future retrieval such as KNN-LM [58], Memorizing Transformer [122], LONGMEM [111], CAMELoT [40] and Memoria [78]. Here CAMELoT and Memoria include some forgetting mechanism to ensure the number of key-values pairs do not grow linearly as the input sequence becomes longer. Most recently, Memory3 [126] proposes to store the knowledge from the pretraining dataset in a knowledge base within the hidden space with $1.1\times 10^{8}$ text chunks with length bounded by 128 tokens. Here the number of text chunks can be easily adjusted and the number of tokens in each chunk can vary thus the size of the memory is flexible.
Memory Module in Text Space. Another line of work proposes to use language models as the abstractor, which inherently extracts the information from the given input and merge them with the existing memory to form a new memory. Here the memory is in the form of a summary of the past seen document. RecurrentGPT [141], as one representative work, proposes to recurrently manage a summarization of the past context. Similarly, MemoryBank [140] designs a hierarchical memory pool, storing (1) the raw conversation history between the user and the agent, (2) the summarization of the conversation, and (3) the summarization of the user’s personalities. Here the last two summaries are in the form of text and serve as the compressed version of the past experiences. Different from RecurrentGPT which is specifically designed for long-context tasks, MemoryBank mainly focuses on conversations. There are also some forgetting mechanisms introduced in MemoryBank to mimic human memory and avoid the ever-growing problem of the memory pool. Moreover, SCM [107] maintains the summary of each conversation and also an interaction embedding to represent the semantics of the interaction.
Memory Module in Symbolic Space. The memory pool in symbolic space includes database as done in ChatDB [41] where the database is updated and queried using the SQL commands generated by the fine-tuned language model. Then Voyager [109] maintains a code base to represent the ever-growing skill library to store and update complex behaviors in Minecraft. These strategies might have the best Long-term Retention and Accuracy Recalling ability as the past experiences are integrated explicitly in the database or the code base. However, real-world experiences encountered over a lifespan may be difficult to encode into a code base (like the Minecraft skill set) or the database. This suggests that while encoding knowledge into symbolic space is suitable for solving specific problems, its applicability in general domains may be limited.

### v2

@@ Abstract
Building a human-like system that continuously interacts with complex environments—whether simulated digital worlds or human society—presents several key challenges. Central to this is enabling continuous, high-frequency interactions, where the interactions are termed experiences. We refer to this envisioned system as the LifeSpan Cognitive System (LSCS). A critical feature of LSCS is its ability to engage in incremental and rapid updates while retaining and accurately recalling past experiences. In this paper we focus on the domain of Large Language Models (LLMs), where we identify two major challenges: (1) Abstraction and Experience Merging, and (2) Long-term Retention with Accurate Recall. These properties are essential for storing new experiences, organizing past experiences, and responding to the environment in ways that leverage relevant historical data. Unlike language models with continual learning, which typically rely on large corpora for fine-tuning and focus on improving performance within specific domains or tasks, LSCS must rapidly and incrementally update with new information from its environment at a high frequency. Existing technologies with the potential of solving the above two major challenges can be classified into four classes based on a conceptual metric called Storage Complexity, which measures the relative space required to store past experiences. Each of these four classes of technologies has its own strengths and limitations while we argue none of them alone can achieve LSCS alone. To this end, we propose a potential instantiation for LSCS that can integrate all four classes of technologies. The new instantiation, serving as a conjecture, operates through two core processes: Absorbing Experiences and Generating Responses.
@@ 1 Introduction
Imagine an environment such as a simulated digital world (Park et al., 2023), a virtual world like Minecraft (Wang et al., 2023b), the Marvel Universe depicted in movies, or even the complexities of human society. Developing an intelligent system capable of engaging with such environments—interacting with its surroundings, absorbing information from experiences, self-evolving based on feedback (Zhou et al., 2024), and living through an entire lifespan, either as a real or virtual entity—remains a significant challenge. When the system inhabits the environment, it interacts with everything around it, including humans, objects, and other elements of the environment. We define these interactions as the system’s experiences, encompassing all of the system’s behaviors and the feedback it receives from the environment—whether visual, physical, linguistic, or across any other modality. This system’s cognitive capabilities include the ability to perceive and interpret its surroundings, retain and recall information through memory, learn and adapt from feedback, and engage in reasoning and decision-making processes. These cognitive functions enable the system to navigate its environment intelligently and evolve over time. We denote the system with the above capabilities as the LifeSpan Cognitive System (LSCS). A real-world example of LSCS is a long-lived, autonomous AI agent—such as a virtual companion or an agent operating within a simulated AI civilization. Over time, it would continuously interact with its environment, accumulate experiences, and refine its internal models. These interactions might span years, requiring the agent to recall past events and lessons learned weeks, months, or years earlier. In this paper, we mainly focus on constructing LSCS based on large language models (LLMs), as LLMs are currently the most promising direction to achieve human-like systems. Moreover, we specifically focus on text domain as most technologies discussed in this paper are constrained in text domain.
@@
Table 1: In this paper, we discuss four categories of technologies with the sub-categories being sub-sections. Here “Abs & ExMg” refers to“Abstraction and Experiences Merging”, "Long-Re & AcRe" means “Long-Term Retention and Accurate Recalling”, and “SC” represents “Storage Complexity”.
@@
To achieve LSCS, some major challenges must be addressed: (1) Abstraction and Experience Merging: The system must distill experiences from its environment by extracting essential information and integrating these abstracted experiences with its existing memories. Interactions with the environment—whether through conversations (Feng et al., 2021; 2022) or visual inputs (Wang et al., 2024h)—often contain redundant information that should be filtered out before storing in memory. Once the key information is abstracted, this crucial information can be used to update the system, merging new abstracted experiences with previous ones while resolving potential conflicts. This process enables the system to acquire new skills, deepen its understanding, correct misconceptions, and continually evolve. (2) Long-Term Retention and Accurate Recalling: For systems that store past experiences in various forms—whether in memory, context, knowledge bases, or model weights—it is crucial to accurately recall relevant information in response to current queries from the environment. This capability necessitates the model’s ability to remember events from the distant past and make informed decisions based on all past experiences. The above two properties distinguish constructing LSCS from existing challenges such as long context problems (Wang et al., 2024d) or continual learning (Wu et al., 2024), where long context problems focus on solving a problem with extremely abundant information and continual learning mainly focuses on enabling models to capture ever-changing world knowledge (Zhang et al., 2023c; Jang et al., 2022a), facilitating self-evolution through human feedback (Tao et al., 2024), or adapting to specific domains (Ke et al., 2023; Yadav et al., 2023).
@@
There are existing technologies that address some of the challenges described above, but each typically falls short of the other. We denote the past experiences as $\mathcal{E}$. Suppose $\mathcal{E}$ can be quantified (such as the number of tokens in the interactions between the system and the environment) and let this quantity be $n$. We then define Storage Complexity as a conceptual measure of the storage requirements beyond model parameters, expressed as a function of $n$, the amount of past experiences. These technologies can be categorized into four groups (illustrated in Figure 1), which will be elaborated in the subsequent sections: Saving $\mathcal{E}$ into Model Parameters (Section 2), Saving $\mathcal{E}$ into Explicit Memory (Section 3), Saving $\mathcal{E}$ into Knowledge Bases for Retrieval (Section 4), and Saving $\mathcal{E}$ into Raw Text and Process All (Section 5). We provide more details of the categorizations in Appendix A. The advantages and potential drawbacks of each approach are discussed in Section 6, with a high-level summary provided in Table 1. This table also outlines the structure of the paper, further subdividing some categories.
@@
With the limitations of existing technologies, we argue that any category of technologies cannot achieve LSCS alone. To deal with this issue, we propose a new possible formulation for LSCS, which integrates all the approaches above. The operational design is depicted in Figure 2, where the process is split into two stages: (1) Abstraction and Merging Experiences when new experiences arise. (2) Generating Responses based on queries from the environment, ensuring long-term retention and precise recall. Each component in this design corresponds to one of the technology categories discussed earlier, and their roles are explored in more detail in Section 7.
@@
Figure 2: The Operating Diagram of LSCS mainly includes two parts: (a) Abstraction and Experiences Merging; (b) Generating responses according to the environmental query, where the ability of long-term retention and accurate recalling should be guaranteed. Note that we add “Notepad” after “Knowledge Base”, which is meant as an analogy. See Section 7.1 for more details.
@@ 2.1 Absorbing $\mathcal{E}$ via Model Editing
Given past experiences, it is possible to construct a knowledge graph from these experiences and then convert the triplets in the graph into factual statements, which could be further injected into the model parameters via model editing (Yao et al., 2023). Here we focus on the methods that directly make edits on the model parameters, i.e., the new knowledge is saved in parameters. Methods such as ROME (Meng et al., 2022), MEMIT (Meng et al., 2023) propose a closed-form solution to edit the MLP layers, while Model-Editing-FT (Gangadhar & Stratos, 2024) fine-tunes the whole model on the factual statements that need to be injected into the model. T-Patcher (Huang et al., 2023c) and CaliNet (Dong et al., 2022) propose to store the new knowledge in additional neurons, however, storing knowledge in additional neurons may inevitably encounter the issue of ever-growing parameters in lifespan settings. As model-editing methods typically focus on storing simple factual statements in the model, they are not directly applicable in the context of LSCS. Thus adjustments are needed for the application.
2.2 Absorbing $\mathcal{E}$ via Continual Learning
@@
Continual learning is usually used in the following situations: (1) Performing real-time assimilation of dynamic data. Existing methods can incorporate information from various sources such as news (Sun et al., 2020; Yu & Ji, 2023), scholarly articles (Cossu et al., 2022), and social media (Cossu et al., 2022). Sun et al. (2020) present ERNIE 2.0, which is a continual pre-training framework that incrementally builds and learns from multiple tasks to maximize knowledge extraction from training data. (2) Injecting a large amount of knowledge into the language model. Related methods using knowledge distillation to preserve existing knowledge is proposed in DER++Buzzega et al. (2020). Jang et al. (2022b) introduce continual knowledge learning, a method for updating temporal knowledge in LLMs, reducing forgetting while acquiring new information. (3) Domain adaptation. Cossu et al. (2022) continually train the model on new data streams for both language and vision, and Ke et al. (2023) introduce a soft-masking mechanism to update language models with domain corpora, aiming to boost performance while preserving general knowledge. Various models are proposed as the domain-adapted version of base language models such as financial domain (Xie et al., 2023), E-commerce domain (Ma et al., 2023), and academic content (Wei et al., 2023). These techniques could all be used for LSCS to adjust to the most up-to-date information. (4) Teaching the system how to speak and how to perform certain types of reasoning. To enable the models to solve novel tasks, task-incremental continual instruction tuning is proposed (Wang et al., 2024e; 2023d). These tasks could be related to mathematical problems (Azerbayev et al., 2023), calculators, search engines, and databases (Qin et al., 2023). With the rapid emergence of new tools like advanced software libraries, novel APIs, or domain-specific utilities  (Liang et al., 2023; Jin et al., 2023), there is a growing need to continually update language models so they can quickly adapt and master these new tools. (5) Adapting to evolving societal values. This leads to continual alignment, where two categories exist: (i) Updating LLMs to align to changing societal values; (ii) Incorporating new demographic groups or value systems into LLMs. Some representative works include CPPO (Zhang et al., 2024b), COPF (Zhang et al., 2023a), COPR (Zhang et al., 2024a), which aim to prevent the forgetting of old policies while injecting incremental preferences objectives into the model. For LSCS, these techniques are important as the system may encounter situations where it needs to be up-to-date, learn plenty of new knowledge, adapt to a new domain if the system enters a new environment, learn a specific kind of tasks, or has to update to meet the societal values.
@@
Catastrophic Forgetting. One major limitation of Saving $\mathcal{E}$ into model parameters is catastrophic forgetting, which is a general problem for methods that require updating model parameters both in model-editing (e.g. in sequential editing scenarios, only LLMs with an external module can work in sequential editing tasks such as WISE (Wang et al., 2024c), GRACE (Hartvigsen et al., 2022) while methods without an external module typically fail (Meng et al., 2022; 2023)) and continual learning. LLMs trained on different stages can encounter unconscious forgetting (Lin et al., 2023), eroding their general knowledge through continual instruction tuning. Studies (Qi et al., 2023) have shown that the behavior of safely aligned LLMs can degrade under these conditions. Some metrics are proposed in TRACE (Wang et al., 2023d) to measure the forgetting of LLMs. While we want to prevent catastrophic forgetting, we do intend to have some extent of forgetting, which is aligned with the ability of Abstraction and Experiences Merging. Different from unlearning part of the knowledge in the LLM which is conducted after having the undesired knowledge in the model (Wang et al., 2024g), we focus on encouraging the model to abstract the information before injecting new experiences into model parameters, where it is acceptable to discard some details. However, Abstracting experiences and discarding unimportant details are not fully explored in continual learning. It is well known that a language model cannot fully remember the training data and instead extracts inherent statistical patterns from the dataset (Carlini et al., 2023). This aligns closely with our goals in LSCS, where we want to abstract events and discard unimportant details. Intuitively, it would be beneficial to control the continual learning process such that only critical information (essential for future decision-making and answering questions) is retained in the model parameters, while less relevant knowledge is discarded. Knowledge Unlearning (Wang et al., 2024g) is related to this objective, as it focuses on removing specific knowledge from the model, which however, may negatively impact model performance (Si et al., 2023). In summary, achieving controllable continual learning is beneficial but under-explored.
@@ 3.1 Methods with a fixed-sized memory module
Most existing works with fixed-sized memory modules have a small memory. Before the transformer era, plenty of works augment recurrent neural networks (RNNs) with a small memory module such as Memory Network (Weston et al., 2014) and its follow-up Sukhbaatar et al. (2015). More advanced methods include TARDIS (Gulcehre et al., 2017), ARMIN (Li et al., 2019), Fast Weights (Ba et al., 2016), Ordered Memory (Shen et al., 2019), etc. Borrowing ideas from human memory, dual-memory system to mimic short-term and long-term memories is proposed (Bidokhti & Ghaemmaghami, 2022a; b). Later with the introduction of transformers, Memory Transformer (Burtsev & Sapunov, 2020) adds memory tokens at the beginning of the sequence to summarize the given sequence, while RMT (Bulatov et al., 2022; 2023) proposes to add memory tokens both at the beginning and the end of the sequence. Here  Bulatov et al. (2023) scales the number of memory tokens to 512 which enables long-term retention of previous information. EMMA (Moro et al., 2023) has a memory pool of similar size, including both short-term and long-term memory. Then Larimar (Das et al., 2024) proposes to use a fixed-sized episodic memory which is shown to be capable of storing up to 1000 factual statements. As opposed to small fixed-sized memory, MemoryLLM (Wang et al., 2024f) introduces a memory for Llama2-7B (Touvron et al., 2023b) encapsulating 1B parameters with 7680 memory tokens per layer, enormously enlarging the memory pool. Differently, Infini-Attention (Munkhdalai et al., 2024) proposes to use linear attention to attend to the past seen tokens, where the stored matrix performs exactly like a fixed-sized memory (of size $d_{key}\times(d_{value}+1)\times H\times L$, where $d_{key},d_{value},H,L$ refer to the dimension of key states, dimension of value states, number of heads and number of layers, respectively) abstracting the past tokens. These fixed-sized memory modules typically come with higher compression than flexible-sized memory modules and also requires much more training for deployment.
@@ 3.2 Methods with flexible-sized memory pool
Memory Module in Hidden Space. Some methods introduce memory slots alongside the inputs to encode information, where the number of memory slots vary according to the length of the input document (Al Adel & Burtsev, 2021). Other techniques store key-value pairs in a memory pool for future retrieval such as KNN-LM (Khandelwal et al., 2019), Memorizing Transformer (Wu et al., 2022), LONGMEM (Wang et al., 2023c), CAMELoT (He et al., 2024b) and Memoria (Park & Bak, 2024). Here CAMELoT and Memoria include some forgetting mechanism to ensure the number of key-values pairs do not grow linearly as the input sequence becomes longer. Most recently, Memory3 (Yang et al., 2024) proposes to store the knowledge from the pretraining dataset in a knowledge base within the hidden space with $1.1\times 10^{8}$ text chunks with length bounded by 128 tokens. Here the number of text chunks can be easily adjusted and the number of tokens in each chunk can vary thus the size of the memory is flexible.
Memory Module in Text Space. Another line of work proposes to use language models as the abstractor, which inherently extracts the information from the given input and merge them with the existing memory to form a new memory. Here the memory is in the form of a summary of the past seen document. RecurrentGPT (Zhou et al., 2023), as one representative work, proposes to recurrently manage a summarization of the past context. Similarly, MemoryBank (Zhong et al., 2023) designs a hierarchical memory pool, storing (1) the raw conversation history between the user and the agent, (2) the summarization of the conversation, and (3) the summarization of the user’s personalities. Here the last two summaries are in the form of text and serve as the compressed version of the past experiences. Different from RecurrentGPT which is specifically designed for long-context tasks, MemoryBank mainly focuses on conversations. There are also some forgetting mechanisms introduced in MemoryBank to mimic human memory and avoid the ever-growing problem of the memory pool. Moreover, SCM (Wang et al., 2023a) maintains the summary of each conversation and also an interaction embedding to represent the semantics of the interaction.
Memory Module in Symbolic Space. The memory pool in symbolic space includes database as done in ChatDB (Hu et al., 2023) where the database is updated and queried using the SQL commands generated by the fine-tuned language model. Then Voyager (Wang et al., 2023b) maintains a code base to represent the ever-growing skill library to store and update complex behaviors in Minecraft. These strategies might have the best Long-term Retention and Accuracy Recalling ability as the past experiences are integrated explicitly in the database or the code base. However, real-world experiences encountered over a lifespan may be difficult to encode into a code base (not like the Minecraft skill set) or the database. This suggests while encoding knowledge into symbolic space could solve specific problems, its may be limited in general domains.